# Sassha: Sharpness-aware Adaptive Second-order Optimization with Stable Hessian Approximation

### v1

@@ 1 Introduction
SGD $265_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 25}}$ $7.29_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.30}}$ $0.703_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.132}}$ $1.31_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.03}}$ $1.260_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.001}}$ $69.32_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.19}}$
Sophia-H $22797_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 10857}}$ $68.15_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 20.19}}$ $8.130_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 3.082}}$ $19.19_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 6.38}}$ $1.463_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.022}}$ $67.76_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.37}}$
AdaHessian $11992_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 5779}}$ $46.94_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 17.60}}$ $4.119_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.136}}$ $12.50_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 6.08}}$ $1.377_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.070}}$ $68.06_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.22}}$
Shampoo $436374_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 9017}}$ $6823.34_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 664.65}}$ $73.27_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 12.51}}$ $49307489_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 56979794}}$ $1.386_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.010}}$ $64.08_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.46}}$
Sassha $\textbf{107}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 40}}$ $\textbf{1.87}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.65}}$ $\textbf{0.238}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.088}}$ $\textbf{0.65}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.86}}$ $\textbf{0.961}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.005}}$ $\textbf{72.14}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.16}}$
@@ 3 Practical Second-order Optimizers Converge to Sharp Minima
We observe that existing second-order optimizers produce solutions with significantly higher sharpness compared to Sassha in all sharpness metrics, which also correlates well with their generalization. We also provide a visualization of the loss landscape for the found solutions, where we find that the solutions obtained by second-order methods are indeed much sharper than that of Sassha (Figure 2).
@@ 4.1 Sharpness-aware Second-order Optimization
$\tilde{f}_{t}(x)\coloneqq f\left(x+\rho\frac{\nabla f(x_{t})}{\|\nabla f(x_{t})\|_{2}}\right),$
@@
$=x_{t}-H\left(x_{t}+\epsilon_{t}^{\star}\right)^{-1}\nabla f\left(x_{t}+\epsilon_{t}^{\star}\right),$
@@
Practical second-order methods must rely on approximately estimated Hessians (i.e., $H\rightarrow\widehat{H}$) since the exact computation is prohibitively expensive for large-scale problems. We choose to employ the diagonal approximation via Hutchinson’s method. However, as we will show in our analysis (Section 6.2), we find that these estimates can become numerically unstable during the sharpness reduction process, as it penalizes Hessian entries close to zero. This can lead to fatal underestimation of the diagonal Hessian compared to scenarios without sharpness minimization, significantly disrupting training. We propose a stable Hessian approximation to address these issue in the following sections.
@@ 4.2 Improving Stability
Approximate second-order methods can yield overly large steps when their diagonal Hessian estimations underestimate the curvature (Dauphin et al., 2015). However, this instability seems to be more present under sharpness minimization, presumably due to smaller top Hessian eigenvalue $\lambda_{1}$ (Agarwala & Dauphin, 2023; Shin et al., 2024) yielding smaller estimated diagonal entries on average:
@@
Conventional techniques such as damping or clipping can be employed to mitigate this, although their additional hyperparameters require careful tuning. Instead, we propose square rooting the Hessian (i.e., $|\widehat{H}|^{1/2}$), which effectively mitigates instability, allowing improved generalization performance over other alternatives without additional hyperparameters. We present empirical validation of this in Section 6.2 and Section E.1.
@@
Absolute Hessian scaling
In neural network training, the computed Hessian often contains negative entries. However, since the square-root operation we introduced applies only to positive values, these entries must first be transformed. To tackle this, we attend to the prior works of Becker et al. (1988); Yao et al. (2021) and employ the absolute function to adjust the negative entries of the diagonal Hessian to be positive, i.e.
@@
where $\widehat{H}_{ii}$ and $\mathbf{e}_{i}$ are the $i^{\text{th}}$ diagonal entry of the approximate diagonal Hessian and the $i^{\text{th}}$ standard basis vector, respectively. Importantly, this also preserve the same optimal rescaling as Newton’s method, which can provide a more effective second-order step compared to alternatives like clipping (Nocedal & Wright, 1999; Murray, 2010; Dauphin et al., 2014; Wang et al., 2013). Additionally, this transformation mitigates the risk of convergence to critical points such as saddle or local maxima. We empirically validate the effectiveness of this approach in Appendix H and Section E.2.
@@ 4.4 Algorithm
4.5 Convergence Analysis
Assumption 4.1.
Assumption 4.2.
Assumption 4.3.
Theorem 4.4.
5 Evaluations
In this section, we demonstrate that Sassha can indeed improve upon existing second-order methods available for standard deep learning tasks. We also show that Sassha performs competitively to the first-order baseline methods. Specifically, Sassha is compared to AdaHessian (Yao et al., 2021), Sophia-H (Liu et al., 2024), Shampoo (Gupta et al., 2018), SGD, AdamW (Loshchilov & Hutter, 2018), and SAM (Foret et al., 2021) on a diverse set of both vision and language tasks. We emphasize that we perform an extensive hyperparameter search to rigorously tune all optimizers and ensure fair comparisons. We provide the details of experiment settings to reproduce our results in Appendix D. The code to reproduce all results reported in this work is made available for download at https://github.com/LOG-postech/Sassha.
5.1 Image Classification
@@
Figure 3: Validation accuracy curves along the training trajectory. We also provide loss curves in Appendix F.
@@
AdamW $175.06$
SAM ${}_{\text{AdamW}}$ $158.06$
AdaHessian $407.69$
Sophia-H $157.60$
Sassha 122.40
@@
We begin by comparing the generalization performance of adaptive second-order methods to that of first-order methods. Across all settings, adaptive second-order methods consistently exhibit lower accuracy than their first-order counterparts. This observation aligns with previous studies indicating that second-order optimization often result in poorer generalization compared to first-order approaches. In contrast, Sassha, benefiting from sharpness minimization, consistently demonstrates superior generalization performance, outperforming both first-order and second-order methods in every setting. Particularly, Sassha is up to 4% more effective than the best-performing adaptive or second-order methods (e.g., WRN-28-10, ViT-s-32). Moreover, Sassha continually surpasses SGD and AdamW, even when they are trained for twice as many epochs, achieving a performance margin of about 0.3% to 3%. Further details are provided in Appendix G.
Interestingly, Sassha also outperforms SAM. Since first-order methods typically exhibit superior generalization performance compared to second-order methods, it might be intuitive to expect SAM to surpass Sassha if the two are viewed merely as the outcomes of applying sharpness minimization to first-order and second-order methods, respectively. However, the results conflict with this intuition. We attribute this to the careful design choices made in Sassha, stabilizing Hessian approximation under sharpness minimization, so as to unleash the potential of the second-order method, leading to its outstanding performance. As a support, we show that naively incorporating SAM into other second-order methods does not yield these favorable results in Appendix H. We also make more comparisons with SAM in Section 5.3.

### v2

@@ 1 Introduction
SGD 265±25.00 7.290±0.300 0.703±0.132 1.310±1.030 1.260±0.001 69.32±0.19
AdaHessian 11992±5779 46.94±17.60 4.119±1.136 12.50±6.080 1.377±0.070 68.06±0.22
Sophia-H 22797±10857 68.15±20.19 8.130±3.082 19.19±6.380 1.463±0.022 67.76±0.37
Shampoo 436374±9017 6823±664.7 73.27±12.51 49307489±56979794 1.386±0.010 64.08±0.46
Sassha 107±40.00 1.870±0.650 0.238±0.088 0.650±0.860 0.961±0.005 72.14±0.16
@@ 3 Practical Second-order Optimizers Converge to Sharp Minima
We observe that existing second-order methods produce solutions with significantly higher sharpness compared to Sassha in all sharpness metrics, which also correlates well with their generalization. We also provide a visualization of the loss landscape for the found solutions, where we find that the solutions obtained by second-order methods are indeed much sharper than that of Sassha (Figure 2).
@@ 4.1 Sharpness-aware Second-order Optimization
$\tilde{f}_{t}(x)\coloneqq f\left(x+\rho\frac{\nabla f(x_{t})}{\|\nabla f(x_{t})\|_{2}}\right),$ (3)
@@
$=x_{t}-H\left(x_{t}+\epsilon_{t}^{\star}\right)^{-1}\nabla f\left(x_{t}+\epsilon_{t}^{\star}\right),$ (4)
@@
Practical second-order methods must rely on approximately estimated Hessians (i.e., $H\rightarrow\widehat{H}$) since the exact computation is prohibitively expensive for large-scale problems. We choose to employ the diagonal approximation via Hutchinson’s method. However, as we will show in our analysis (Section 6.2), we find that these estimates can become numerically unstable during the sharpness reduction process, as it penalizes Hessian entries close to zero. This can lead to fatal underestimation of the diagonal Hessian compared to scenarios without sharpness minimization, significantly disrupting training. We propose a stable Hessian approximation to address these issues in the following sections.
@@ 4.2 Improving Stability
Approximate second-order methods can yield overly large steps when their diagonal Hessian estimations underestimate the curvature (Dauphin et al., 2015). However, this instability seems to be more present under sharpness minimization, presumably due to smaller top Hessian eigenvalue $\lambda_{1}$ (Agarwala & Dauphin, 2023; Shin et al., 2025) yielding smaller estimated diagonal entries on average:
@@
A conventional approach to mitigating this involves damping or clipping, which, while stabilizing it reasonably well, requires carefully tuning their additional hyperparameters. Instead, we propose square rooting the Hessian (i.e., $|\widehat{H}|^{1/2}$), which effectively mitigates instability and allows improved generalization performance over other alternatives without additional hyperparameters. We present empirical validation of this in Section 6.2 and Section F.1.
@@
Second, it can further be interpreted as a geometric interpolation between the identity matrix $I$ and the preconditioning matrix $|\widehat{H}|^{\alpha}$, which, as theoretically analyzed in (Amari et al., 2021), provide a natural mechanism for balancing between bias and variance of the population risk, thereby improving generalization. We specifically adopt $\alpha=1/2$ (i.e., square root), as it has consistently demonstrated robust performance across various scenarios (Amari et al., 2021; Kingma & Ba, 2015).
Avoiding critical points.
A well-known limitation of applying second-order optimization to deep learning objectives is the risk of convergence to saddle points or local maxima. To mitigate this, we attend to the prior works of Becker et al. (1988); Yao et al. (2021) and employ the absolute function to conservatively adjust the negative entries of the diagonal Hessian to be positive, i.e.
@@
where $\widehat{H}_{ii}$ and $\mathbf{e}_{i}$ are the $i^{\text{th}}$ diagonal entry of the approximate diagonal Hessian and the $i^{\text{th}}$ standard basis vector, respectively. Here, the basic idea is to maintain the directionality of the gradient by flipping the sign of the negative entries in the Hessian, preserving the original magnitude of its values. This allows our method not only to take descent steps along directions of originally negative curvature, but also to faithfully preserve the relative scale specifically among the negative elements of the Hessian. As a result, it mitigates the risk of convergence to the undesired critical points. We empirically validate the effectiveness of this approach in Section F.2.
@@ 4.4 Algorithm
4.5 Convergence Analysis
Assumption 4.1.
Assumption 4.2.
Assumption 4.3.
Theorem 4.4.
@@ 4.5 Convergence Analysis
4.6 Flatness Analysis
To understand how Sassha can end up in flatter minima as observed in Section 3, we attend to linear stability analysis. Originally developed to explain similar behavior in SGD (Wu et al., 2018) and also later extended to SAM (Shin et al., 2025), this framework suggests that an optimizer does not settle in every minimum it approaches, but instead escapes unless it encounters one that satisfies specific stability conditions—conditions that can vary between optimizers. Based on this, we demonstrate that Sassha requires a minimum to possess a certain level of flatness and Hessian uniformity to settle in it, whereas approximate second-order optimizers do not necessarily require such restriction, thus allowing it to stay in much sharper minima.
Consider a general optimizer $x_{t+1}=x_{t}-G(x_{t};\xi_{t})$ with randomness induced by i.i.d. variables $\xi_{t}$ that are independent from the iterate $x_{t}$. Also, let us assume that the minima possess a fixed point $x^{\star}$ such that $\nabla f_{\xi}(x^{\star})=0$ for any $\xi$.
With these, we define the linear stability of the fixed point $x^{\star}$ as follows:
Definition 4.5.
(Linear stability). A fixed point $x^{\star}$ is linearly stable for the optimizer $G$ if there exists a constant $C$ such that
$\mathbb{E}\bigl{[}\lVert\hat{x}_{t}-x^{\star}\rVert^{2}\bigr{]}\leq C\lVert\hat{x}_{0}-x^{\star}\rVert^{2},\text{ for all }t>0$
under the linearized dynamics near $x^{\star}$: $\hat{x}_{t+1}=\hat{x}_{t}-\nabla G(x^{\star})(\hat{x}_{t}-x^{\star})$, i.e., when it does not deviate infinitely far from $x^{\star}$.
Here, this linear dynamic arises when $x_{t}$ has approached sufficiently close to $x^{\star}$ such that the loss becomes approximately quadratic.
Under this framework, we present the necessary conditions of the linearly stable minima of Sassha in the following corollary.
Corollary 4.6.
Assume without loss of generality that $x^{\star}=0$. Then, the linearly stable fixed point $x^{\star}$ of Sassha satisfy the following necessary conditions:
$0\leq$ $a(1+\rho a)\leq\frac{2\epsilon}{\eta},\quad 0\leq s^{2}_{2}\leq\frac{\epsilon^{2}}{\eta^{2}-2\eta\rho\epsilon},$
$0\leq$ $s^{2}_{3}\leq\frac{\epsilon^{2}}{2\eta^{2}\rho},\quad 0\leq s^{2}_{4}\leq\frac{\epsilon^{2}}{\eta^{2}\rho^{2}},$ (6)
where $a=\lambda_{max}(\mathbb{E}[H_{\xi}])$ and $s_{k}=\lambda_{max}((\mathbb{E}[H_{\xi}^{k}]-\mathbb{E}[H_{\xi}]^{k})^{1/k})$ are the sharpness and the non-uniformity of the stochastic Hessian measured with the $k$-th moment, respectively.
A detailed proof is provided in Appendix D.
These results indicate that Sassha escapes from minima unless they satisfies both low sharpness and uniformly distributed Hessian moments, with the conditions becoming much stricter with larger $\rho$ and $\eta$. In comparison, standard approximate second-order methods of the form $x_{t+1}=x_{t}-\eta P_{t}\nabla f(x_{t})$ with $P\approx H^{-1}$ remain stable without such conditions, as shown by Wu et al. (2018):
$\lambda_{\max}(P^{-1}H)\approx 1\leq\frac{2}{\eta},$
allowing convergence to minima of any sharpness provided $\eta\leq 2$.
@@ 4.6 Flatness Analysis
Figure 3: Validation accuracy curves along the training trajectory. We also provide loss curves in Appendix G.
5 Evaluations
In this section, we demonstrate that Sassha can indeed improve upon existing second-order methods available for standard deep learning tasks. We also show that Sassha performs competitively to the first-order baseline methods. Specifically, Sassha is compared to AdaHessian (Yao et al., 2021), Sophia-H (Liu et al., 2024), Shampoo (Gupta et al., 2018), SGD, AdamW (Loshchilov & Hutter, 2018), and SAM (Foret et al., 2021) on a diverse set of both vision and language tasks. We emphasize that we perform an extensive hyperparameter search to rigorously tune all optimizers and ensure fair comparisons. We provide the details of experiment settings to reproduce our results in Appendix E. The code to reproduce all results reported in this work is made available for download at https://github.com/LOG-postech/Sassha.
@@ 5 Evaluations
AdamW $175.06_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.19}}$
SAM ${}_{\text{AdamW}}$ $158.06_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.23}}$
AdaHessian $407.69_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.20}}$
Sophia-H $157.60_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.37}}$
Sassha $\textbf{122.40}_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.16}}$
@@ 5.1 Image Classification
We begin by comparing the generalization performance of adaptive second-order methods to that of first-order methods. Across all settings, adaptive second-order methods consistently exhibit lower accuracy than their first-order counterparts. This observation aligns with previous studies indicating that second-order optimization often result in poorer generalization compared to first-order approaches. In contrast, Sassha, benefiting from sharpness minimization, consistently demonstrates superior generalization performance, outperforming both first-order and second-order methods in every setting. Particularly, Sassha is up to 4% more effective than the best-performing adaptive or second-order methods (e.g., WRN-28-10, ViT-s-32). In addition, Sassha continually surpasses SGD and AdamW by approximately 0.3% to 3%, even when these methods are trained for twice as many epochs. Further details on these experiments are provided in Appendix H.
Interestingly, Sassha also outperforms SAM. Since first-order methods typically exhibit superior generalization performance compared to second-order methods, it might be intuitive to expect SAM to surpass Sassha if the two are viewed merely as the outcomes of applying sharpness minimization to first-order and second-order methods, respectively. However, the results conflict with this intuition. We attribute this to the careful design choices made in Sassha, stabilizing Hessian approximation under sharpness minimization, so as to unleash the potential of the second-order method, leading to its outstanding performance. As a support, we show that naively incorporating SAM into other second-order methods does not yield these favorable results in Appendix I. We also make more comparisons with SAM in Section 5.3.